OpenStudy (anonymous):

What is the sum or difference? 5x8 + 9x8 (1 point) 45x8 –4x8 14x16 14x8

OpenStudy (raden):

use distributive property : a*b + c*b = (a+c)b

OpenStudy (anonymous):

Take 8 common..

OpenStudy (raden):

5x8 + 9x8 = (5+9)8 = 14*8